tyson:

Solve for x: 7x + 15 = 38

Hero:

@tyson begin by subtracting 15 from both sides.

AngeI:

Ok so for this one you have to subtract 15 from both sides 7x+15(-15)=38(-15) Then you must divide what is left by 7 to find what x is

Hero:



Hero:

@tyson if you click the little pencil icon, you can finish the work

tyson:

i did all that i just need to know the anwser to the problem i couldnt figure out the division part

Hero:

In that case, you should show the work you've done up to that point.

tyson:

i wont let me i have been tryiing

tyson:

nvm gtg bye

russkaya05:

Let's solve your equation step-by-step. 7x+15=38 Step 1: Subtract 15 from both sides. 7x+15−15=38−15 7x=23 Step 2: Divide both sides by 7. 7\[7x/7\] = \[23/7\]

russkaya05:

opps

russkaya05:

3 = 23/7

tyson:

back

russkaya05:

what?